### v1

@@ 3.4 Inference via cluster jackknifing
Because the model-robust standardization estimator can accommodate different specifications of the working outcome regression models, it is desirable to have a one-size-fits-all solution to statistical inference. Therefore, we consider a deletion-based cluster jackknife variance estimator that is simple to implement. The jackknife variance estimator has been previously studied for model-based inference with a small number of clusters. For example, in the context of linear regression of clustered data, many have shown that jackknife variance estimator often outperforms the bias-corrected cluster-robust variance estimators in terms of the test size in small samples.46 The jackknife variance estimator is also connected with the Mancl and DeRouen GEE variance estimator to combat finite-sample bias.47, 48 Finally, compared to other resampling-based methods such as cluster bootstrap and permutation, the computation cost needed for jackknife is more affordable as one only needs to refit the model-robust standardization estimator at most $m$ times.
@@
Hypothesis tests and confidence intervals can then be constructed with the $t(m-1)$ approximate sampling distribution.
@@ 3.5 A model-robust test for informative cluster size
To proceed, when $f$ is the difference measure, we first compute the difference between the cluster-ATE and individual-ATE estimators as $\widehat{D}=\widehat{\Delta}_{C}-\widehat{\Delta}_{I}=(\widehat{\mu}_{C}(1)-\widehat{\mu}_{C}(0))-(\widehat{\mu}_{I}(1)-\widehat{\mu}_{I}(0))$. To estimate the variance of $\widehat{D}$, we can use the jackknife resampling method described in Section 3.4. Specifically, we sequentially leave out one cluster at a time and recompute the difference estimators on the remaining $m-1$ clusters to obtain leave-one-cluster-out estimators $\widehat{D}^{(-i)}=\widehat{\Delta}_{C}^{(-i)}-\widehat{\Delta}_{I}^{(-i)}$ for each $i=1,\ldots,m$. The jackknife estimate of the variance of $\widehat{D}$ is then given by
@@
where $\overline{D}=\frac{1}{m}\sum_{i=1}^{m}\widehat{D}^{(-i)}$ is the average of the jackknife estimators. The final test statistic is constructed as $T=\widehat{D}/\sqrt{\widehat{V}_{D}}$, which is approximately $t$-distributed with $m-1$ degrees of freedom under the null hypothesis. This consists of the basis for computing p-values. The procedures for carrying out the test when $f$ defines the ratio effect follow a similar construction, except that one may define the test based on $\log(\widehat{\Delta}_{C})-\log(\widehat{\Delta}_{I})$ to improve numerical stability.
@@ 4 Simulation experiments
We conduct a series of simulation studies to compare model-robust standardization to standard practice of estimating regression coefficient. Throughout, we fix the number of simulations to $1000$ and focus on estimating $\Delta_{C}$ and $\Delta_{I}$; the choice of effect measure and $f$ depends on the type of the outcome and will be explained in each setting. In Sections 4.1-4.4, we investigate the 4 combinations arising from 2 outcome types (continuous or binary) and 2 scenarios for informative cluster sizes (absence or presence). Tables 1-4 summarize the main results, and the corresponding Monte Carlo standard errors quantifying simulation uncertainty are given in Web Tables 6-9; the definitions of each uncertainty measure are given in Table 6 of Morris et al.49 Section 4.5 summarizes findings from additional simulations presented in the Web Appendix and Section 4.6 illustrates the performance of the test for informative cluster size.
@@ 4.1 Non-informative cluster size and a continuous outcome
The random intercept $\gamma_{i}\sim\mathcal{N}(0,0.2)$ and induces a positive adjusted ICC $\approx 0.17$ under the treatment condition. In this case, we define the target estimands on the difference scale ($f(x,y)=x-y$), and analytically derive the cluster-ATE and individual-ATE as $-3$. We draw $A_{i}\sim\text{Bernoulli}(0.5)$, and set the observed outcome $Y_{ij}=A_{i}Y_{ij}(1)+(1-A_{i})Y_{ij}(0)$. Here, the true potential outcome model assumes a constant treatment effect across individuals but incorporates relatively nonlinear covariate functional forms. This design allows us to evaluate the performance of the model-robust standardization and model-based coefficient estimators under model misspecification, where the fitted models are allowed to differ from the true data generating mechanisms. Second, the cluster size $N_{i}$ is neither associated with $Y_{ij}(a)$ nor acts as an effect modifier. Consequently, there is no informative cluster size, ensuring equivalence between the cluster-ATE and the individual-ATE.
Table 1: Simulation results under a data generating process with non-informative cluster size and a continuous outcome. LM: linear model based on cluster-level means; LMM: linear mixed model based on individual-level data; GEE-exch: marginal model fitted with GEE and an exchangeable working correlation structure; GEE-ind: marginal model fitted with GEE and working independence. Monte Carlo standard errors for performance metric estimates are presented in Web Table 1.
$m$ Covariate Outcome model Method BIAS MCSD AESE COV BIAS MCSD AESE COV
adjustment
@@
[…]
MRS -0.6 1.72 1.73 94.7 -0.3 1.84 1.90 93.8
[…]
MRS -0.6 1.72 1.73 94.7 -0.3 1.84 1.90 93.8
[…]
MRS -0.6 1.72 1.73 94.7 -0.3 1.84 1.90 93.8
[…]
MRS -0.4 1.74 1.74 94.7 -0.1 1.85 1.92 93.9
[…]
MRS 0.2 0.30 0.38 98.3 0.2 0.32 0.40 97.6
[…]
MRS 0.2 0.30 0.35 97.2 0.2 0.32 0.36 96.7
[…]
MRS 0.2 0.30 0.35 97.2 0.2 0.32 0.36 96.7
[…]
MRS 0.0 0.30 0.36 97.0 0.2 0.32 0.37 96.6
@@
[…]
MRS -0.2 0.49 0.48 95.3 0.0 0.54 0.53 94.2
[…]
MRS -0.2 0.49 0.48 95.3 0.0 0.54 0.53 94.2
[…]
MRS -0.2 0.49 0.48 95.3 0.0 0.54 0.53 94.2
[…]
MRS -0.2 0.49 0.48 95.5 0.0 0.54 0.53 94.3
[…]
MRS 0.0 0.17 0.18 96.9 0.0 0.18 0.19 96.1
[…]
MRS 0.0 0.17 0.18 96.5 0.0 0.18 0.19 95.8
[…]
MRS 0.0 0.17 0.18 96.4 0.0 0.18 0.19 95.8
[…]
MRS 0.0 0.17 0.18 96.3 0.0 0.18 0.19 96.2
@@
We examine $4$ working outcome models. These are (W1) a linear regression model based on cluster-level means described in Example 3.1; (W2) a linear mixed model based on individual-level observations described in Example 3.2; (W3) a marginal model fitted with GEE under the exchangeable working correlation structure; and (W4) a marginal model fitted with GEE under working independence, as described in Example 3.4. For each (W1)-(W4), we consider both an unadjusted version where only the treatment indicator $A_{i}$ is included, and a covariate-adjusted version where additional baseline characteristics, $\{X_{1ij},X_{2ij},H_{1i},H_{2i},N_{i}\}$ are controlled for in the regression as linear main effects. For each one of these $8$ working models, we then obtain the model-robust standardization (MRS) estimator as well as the model-based coefficient (Coef) estimator corresponding to the treatment variable $A_{i}$. The standard errors for the MRS estimators are computed via cluster jackknifing, whereas those for the Coef estimators are obtained from the heteroskedasticity-robust sandwich variance estimator (for (W1) based on the sandwich R package), the approximate jackknife variance estimator (for (W2) based on the clubSandwich R package), and the Mancl and DeRouen bias-corrected variance estimator (for (W3) and (W4) based on the geesmv R package). These choices of standard error estimators reflect existing recommendations for finite-sample adjustments in CRTs.47, 50, 51, 48 We investigate four performance metrics including the percentage bias (BIAS), Monte Carlo standard deviation (MCSD), average estimated standard error (AESE), and coverage probabilities for the 95% confidence intervals constructed with the $t(m-1)$ distribution. The performance metrics are defined relative to the estimand $\Delta_{C}=\Delta_{I}=-3$.
@@
Table 1 presents the simulation results under non-informative cluster size for a continuous outcome. We observe that the Coef estimators have negligible bias to either the cluster-ATE and individual-ATE, a result that is consistent with prior studies.11, 22 The MRS estimators also have no larger bias than Coef estimators, regardless of the choice of regression models. Second, the variance estimator for Coef tends to underestimate the true variance, resulting in approximate or below 90% coverage, when $m=30$ and covariate-adjusted working models (W1)-(W2) are used. In contrast, the MRS estimators, at worst, provide conservative coverage under comparable scenarios. Third, covariate adjustment can improve efficiency regardless of sample size and choice of estimation method. To summarize, under non-informative cluster size, the MRS method consistently performs no worse than the Coef method with a continuous outcome, although the latter has little bias despite model misspecification. This simulation experiment paves the way for Section 4.2 where we modify the data generating mechanism to induce informative cluster size with a continuous outcome.
@@ 4.2 Informative cluster size and a continuous outcome
Table 2: Simulation results under a data generating process with informative cluster size and a continuous outcome. LM: linear model based on cluster-level means; LMM: linear mixed model based on individual-level data; GEE-exch: marginal model fitted with GEE and an exchangeable working correlation structure; GEE-ind: marginal model fitted with GEE and working independence. Monte Carlo standard errors for performance metric estimates are presented in Web Table 2.
$m$ Covariate Outcome model Method BIAS MCSD AESE COV BIAS MCSD AESE COV
adjustment
@@
$30$$\times$ (W1) LM Coef –0.3 2.16 2.12 93.6 –27.6 2.16 2.12 78.5
MRS –0.3 2.16 2.25 94.5 –2.1 2.77 2.88 93.7
(W2) LMM Coef –0.2 2.16 2.12 93.6 –27.6 2.16 2.12 78.5
MRS –0.3 2.16 2.25 94.5 –2.1 2.77 2.88 93.7
(W3) GEE-exch Coef -0.3 2.16 2.28 94.6 -27.7 2.16 2.28 81.7
MRS –0.3 2.16 2.25 94.5 –2.1 2.77 2.88 93.7
(W4) GEE-ind Coef 35.0 2.80 2.99 92.7 -2.0 2.80 2.99 93.8
MRS -0.3 2.18 2.31 94.5 -2.0 2.80 2.94 93.4
$\checkmark$ (W1) LM Coef 0.1 1.18 1.04 90.7 -27.3 1.18 1.04 47.9
MRS 0.1 1.18 1.55 97.4 -7.3 1.28 1.71 94.5
(W2) LMM Coef 0.2 1.18 1.04 90.7 -27.3 1.18 1.04 48.0
MRS 0.1 1.18 1.43 96.5 -7.3 1.28 1.58 93.6
(W3) GEE-exch Coef 0.1 1.18 1.46 96.8 -27.3 1.18 1.46 68.7
MRS 0.1 1.18 1.43 96.5 -7.3 1.28 1.58 93.6
(W4) GEE-ind Coef 29.9 1.31 1.66 83.8 -5.7 1.31 1.66 94.7
MRS 1.6 1.20 1.49 97.0 -5.7 1.31 1.63 94.1
@@
$100$$\times$ (W1) LM Coef –0.2 0.72 0.74 95.9 –28.4 0.72 0.74 34.6
MRS –0.2 0.72 0.75 96.3 –1.1 0.85 0.89 95.2
(W2) LMM Coef 0.0 0.72 0.74 95.8 –28.3 0.72 0.74 35.3
MRS –0.2 0.72 0.75 96.3 –1.1 0.85 0.89 95.2
(W3) GEE-exch Coef -0.3 0.72 0.75 96.3 -28.5 0.72 0.75 35.9
MRS –0.2 0.72 0.75 96.3 –1.1 0.85 0.89 95.2
(W4) GEE-ind Coef 37.9 0.83 0.87 51.2 -1.0 0.83 0.87 95.5
MRS -0.2 0.74 0.76 96.2 -1.0 0.83 0.86 95.4
$\checkmark$ (W1) LM Coef 0.2 0.47 0.44 92.9 -28.1 0.47 0.44 3.7
MRS 0.2 0.47 0.49 95.0 -2.1 0.52 0.54 93.8
(W2) LMM Coef 0.7 0.47 0.44 93.1 -27.8 0.47 0.44 3.7
MRS 0.2 0.47 0.48 94.9 -2.1 0.52 0.52 93.4
(W3) GEE-exch Coef 0.3 0.47 0.48 94.9 -28.0 0.47 0.48 4.9
MRS 0.2 0.47 0.48 94.9 -2.1 0.52 0.53 93.4
(W4) GEE-ind Coef 37.1 0.51 0.53 10.8 -1.6 0.51 0.53 93.6
MRS 0.6 0.47 0.49 94.7 -1.6 0.51 0.52 93.4
@@
Table 2 presents the simulation results when the exact same 4 working models (with and without linear covariate adjustment) in Section 4.1 are fitted; the presentation of Table 2 follows Table 1 to enable direct comparisons. It is apparent that, under informative cluster size, the performance of the Coef estimator critically depends on the choice of the outcome regression model. In particular, regardless of covariate adjustment, the Coef estimator from the independence GEE (W4) is biased for estimating cluster-ATE, while the Coef estimators from the remaining models, (W1)-(W3), are biased for estimating individual-ATE. This bias then becomes the primary source for under-coverage. Such biases have been pointed out from previous studies,11, 13 and arise because the fitted regression model does not employ the right weighting scheme that matches the marginal estimands $\Delta_{C}$ or $\Delta_{I}$ (also see Kahan et al.5 and Lee et al.12 for detailed explanation about why Coef from LMM could be biased under informative cluster size). In sharp contrast, the relative bias is negligible and the empirical coverage probabilities are close to $95\%$ for MRS estimators, regardless of the choice of outcome regression models and covariate adjustment. Finally, covariate adjustment under MRS also leads to no worse finite-sample efficiency compared to no covariate adjustment. In summary, although Coef and MRS perform relatively similarly under non-informative cluster size, the findings in Section 4.2 suggest that (1) the performance of Coef highly depends on the choice of outcome model and Coef may target neither cluster-ATE or individual-ATE; (2) even with a given outcome model whose Coef estimator is biased for the estimand of interest, applying our MRS procedure can fix the bias.

### v2

@@ 3.4 Inference via cluster jackknifing
Because the model-robust standardization estimator can accommodate different specifications of the working outcome regression models, it is desirable to have a one-size-fits-all solution to statistical inference. Therefore, we consider a deletion-based cluster jackknife variance estimator that is simple to implement. The jackknife variance estimator has been previously studied for model-based inference with a small number of clusters. For example, in the context of linear regression of clustered data, many have shown that jackknife variance estimator often outperforms the bias-corrected cluster-robust variance estimators in terms of the test size in small samples.45 The jackknife variance estimator is also connected with the Mancl and DeRouen GEE variance estimator to combat finite-sample bias.46, 47 Finally, compared to other resampling-based methods such as cluster bootstrap and permutation, the computation cost needed for jackknife is more affordable as one only needs to refit the model-robust standardization estimator at most $m$ times.
@@
Hypothesis tests and confidence intervals can then be constructed with the t(m1)t(m-1) approximate sampling distribution.
@@ 3.5 A model-robust test for informative cluster size
To proceed, when $f$ is the difference measure, we first compute the difference between the cluster-ATE and individual-ATE estimators as $\widehat{D}=\widehat{\Delta}_{C}-\widehat{\Delta}_{I}=(\widehat{\mu}_{C}(1)-\widehat{\mu}_{C}(0))-(\widehat{\mu}_{I}(1)-\widehat{\mu}_{I}(0))$. To estimate the variance of $\widehat{D}$, we can use the jackknife resampling method described in Section 3.4. Specifically, we sequentially leave out one cluster at a time and recompute the difference estimators on the remaining m1m-1 clusters to obtain leave-one-cluster-out estimators $\widehat{D}^{(-i)}=\widehat{\Delta}_{C}^{(-i)}-\widehat{\Delta}_{I}^{(-i)}$ for each $i=1,\ldots,m$. The jackknife estimate of the variance of $\widehat{D}$ is then given by
@@
where $\overline{D}=\frac{1}{m}\sum_{i=1}^{m}\widehat{D}^{(-i)}$ is the average of the jackknife estimators. The final test statistic is constructed as $T=\widehat{D}/\sqrt{\widehat{V}_{D}}$, which is approximately $t$-distributed with m1m-1 degrees of freedom under the null hypothesis. This consists of the basis for computing p-values. The procedures for carrying out the test when $f$ defines the ratio effect follow a similar construction, except that one may define the test based on $\log(\widehat{\Delta}_{C})-\log(\widehat{\Delta}_{I})$ to improve numerical stability.
@@ 4 Simulation experiments
We conduct a series of simulation studies to compare model-robust standardization to standard practice of estimating regression coefficient. Throughout, we fix the number of simulations to $1000$ and focus on estimating $\Delta_{C}$ and $\Delta_{I}$; the choice of effect measure and $f$ depends on the type of the outcome and will be explained in each setting. In Sections 4.1-4.4, we investigate the 4 combinations arising from 2 outcome types (continuous or binary) and 2 scenarios for informative cluster sizes (absence or presence). Tables 1-4 summarize the main results, and the corresponding Monte Carlo standard errors quantifying simulation uncertainty are given in Web Tables 1-4; the definitions of each uncertainty measure are given in Table 6 of Morris et al.48 Section 4.5 summarizes findings from additional simulations presented in the Web Appendix and Section 4.6 illustrates the performance of the test for informative cluster size.
@@ 4.1 Non-informative cluster size and a continuous outcome
The random intercept $\gamma_{i}\sim\mathcal{N}(0,0.2)$ and induces a positive adjusted ICC $\approx 0.17$ under the treatment condition. In this case, we define the target estimands on the difference scale (f(x,y)=xyf(x,y)=x-y), and analytically derive the cluster-ATE and individual-ATE as 3-3. We draw $A_{i}\sim\text{Bernoulli}(0.5)$, and set the observed outcome $Y_{ij}=A_{i}Y_{ij}(1)+(1-A_{i})Y_{ij}(0)$. Here, the true potential outcome model assumes a constant treatment effect across individuals but incorporates relatively nonlinear covariate functional forms. This design allows us to evaluate the performance of the model-robust standardization and model-based coefficient estimators under model misspecification, where the fitted models are allowed to differ from the true data generating mechanisms. Second, the cluster size $N_{i}$ is neither associated with $Y_{ij}(a)$ nor acts as an effect modifier. Consequently, there is no informative cluster size, ensuring equivalence between the cluster-ATE and the individual-ATE.
Table 1: Simulation results under a data generating process with non-informative cluster size and a continuous outcome. LM: linear model based on cluster-level means; LMM: linear mixed model based on individual-level data; GEE-exch: marginal model fitted with GEE and an exchangeable working correlation structure; GEE-ind: marginal model fitted with GEE and working independence. Each empirical coverage probability is reported together with its Monte Carlo 95% confidence limits (95% CL). Additional Monte Carlo standard errors for performance metric estimates are presented in Web Table 1.
$m$ Covariate Outcome model Method BIAS MCSD COV BIAS MCSD COV
adjustment (AESE) (95% CL) (AESE) (95% CL)
@@
$30$ $\times$ (W1) LM Coef -0.6 1.72 (1.63) 93.2 (91.8, 94.6) \ \ \
MRS -0.6 1.72 (1.73) 94.7 (93.5, 95.9) -0.3 1.84 (1.90) 93.8 (92.5, 95.1)
(W2) LMM Coef -0.6 1.72 (1.63) 93.2 (91.8, 94.6) \ \ \
MRS -0.6 1.72 (1.73) 94.7 (93.5, 95.9) -0.3 1.84 (1.90) 93.8 (92.5, 95.1)
(W3) GEE-exch Coef -0.6 1.72 (1.76) 95.1 (94.0, 96.2) \ \ \
MRS -0.6 1.72 (1.73) 94.7 (93.5, 95.9) -0.3 1.84 (1.90) 93.8 (92.5, 95.1)
(W4) GEE-ind Coef -0.1 1.85 (1.95) 94.4 (93.2, 95.6) \ \ \
MRS -0.4 1.74 (1.74) 94.7 (93.5, 95.9) -0.1 1.85 (1.92) 93.9 (92.6, 95.2)
$\checkmark\checkmark$ (W1) LM Coef 0.2 0.30 (0.25) 90.7 (89.1, 92.3) \ \ \
MRS 0.2 0.30 (0.38) 98.3 (97.5, 99.1) 0.2 0.32 (0.40) 97.6 (96.6, 98.6)
(W2) LMM Coef 0.2 0.30 (0.25) 90.8 (89.2, 92.4) \ \ \
MRS 0.2 0.30 (0.35) 97.2 (96.2, 98.2) 0.2 0.32 (0.36) 96.7 (95.6, 97.8)
(W3) GEE-exch Coef 0.2 0.30 (0.35) 97.6 (96.7, 98.5) \ \ \
MRS 0.2 0.30 (0.35) 97.2 (96.2, 98.2) 0.2 0.32 (0.36) 96.7 (95.6, 97.8)
(W4) GEE-ind Coef 0.2 0.32 (0.38) 96.7 (95.7, 97.7) \ \ \
MRS 0.0 0.30 (0.36) 97.0 (96.0, 98.0) 0.2 0.32 (0.37) 96.6 (95.5, 97.7)
@@
$100$ $\times$ (W1) LM Coef -0.2 0.49 (0.47) 94.8 (93.7, 95.9) \ \ \
MRS -0.2 0.49 (0.48) 95.3 (94.2, 96.4) 0.0 0.54 (0.53) 94.2 (93.0, 95.4)
(W2) LMM Coef -0.2 0.49 (0.47) 94.8 (93.7, 95.9) \ \ \
MRS -0.2 0.49 (0.48) 95.3 (94.2, 96.4) 0.0 0.54 (0.53) 94.2 (93.0, 95.4)
(W3) GEE-exch Coef -0.2 0.49 (0.48) 95.5 (94.5, 96.5) \ \ \
MRS -0.2 0.49 (0.48) 95.3 (94.2, 96.4) 0.0 0.54 (0.53) 94.2 (93.0, 95.4)
(W4) GEE-ind Coef 0.0 0.54 (0.54) 94.6 (93.5, 95.7) \ \ \
MRS -0.2 0.49 (0.48) 95.5 (94.5, 96.5) 0.0 0.54 (0.53) 94.3 (93.1, 95.5)
$\checkmark\checkmark$ (W1) LM Coef 0.0 0.17 (0.16) 94.2 (93.1, 95.3) \ \ \
MRS 0.0 0.17 (0.18) 96.9 (96.0, 97.8) 0.0 0.18 (0.19) 96.1 (95.1, 97.1)
(W2) LMM Coef 0.0 0.17 (0.16) 93.9 (92.8, 95.0) \ \ \
MRS 0.0 0.17 (0.18) 96.5 (95.6, 97.4) 0.0 0.18 (0.19) 95.8 (94.8, 96.8)
(W3) GEE-exch Coef 0.0 0.17 (0.18) 96.6 (95.7, 97.5) \ \ \
MRS 0.0 0.17 (0.18) 96.4 (95.5, 97.3) 0.0 0.18 (0.19) 95.8 (94.8, 96.8)
(W4) GEE-ind Coef 0.0 0.18 (0.19) 96.2 (95.3, 97.1) \ \ \
MRS 0.0 0.17 (0.18) 96.3 (95.4, 97.2) 0.0 0.18 (0.19) 96.2 (95.3, 97.1)
@@
We examine $4$ working outcome models. These are (W1) a linear regression model based on cluster-level means described in Example 3.1; (W2) a linear mixed model based on individual-level observations described in Example 3.2; (W3) a marginal model fitted with GEE under the exchangeable working correlation structure; and (W4) a marginal model fitted with GEE under working independence, as described in Example 3.4. For each (W1)-(W4), we consider both an unadjusted version where only the treatment indicator $A_{i}$ is included, and a covariate-adjusted version where additional baseline characteristics, $\{X_{1ij},X_{2ij},H_{1i},H_{2i},N_{i}\}$ are controlled for in the regression as linear main effects. For each one of these $8$ working models, we then obtain the model-robust standardization (MRS) estimator as well as the model-based coefficient (Coef) estimator corresponding to the treatment variable $A_{i}$. The standard errors for the MRS estimators are computed via cluster jackknifing, whereas those for the Coef estimators are obtained from the heteroskedasticity-robust sandwich variance estimator (for (W1) based on the sandwich R package), the approximate jackknife variance estimator (for (W2) based on the clubSandwich R package), and the Mancl and DeRouen bias-corrected variance estimator (for (W3) and (W4) based on the geesmv R package). These choices of standard error estimators reflect existing recommendations for finite-sample adjustments in CRTs.46, 49, 47 We investigate four performance metrics including the percentage bias (BIAS), Monte Carlo standard deviation (MCSD), average estimated standard error (AESE), and coverage probabilities for the 95% confidence intervals constructed with the t(m1)t(m-1) distribution ($95\%$ confidence limits of empirical coverage48 also provided to quantify uncertainty). The performance metrics are defined relative to the estimand $\Delta_{C}=\Delta_{I}=-3$.
@@
Table 1 presents the simulation results under non-informative cluster size for a continuous outcome. We observe that the Coef estimators have negligible bias to either the cluster-ATE and individual-ATE, a result that is consistent with prior studies.11, 21 The MRS estimators also have no larger bias than Coef estimators, regardless of the choice of regression models. Second, the variance estimator for Coef tends to underestimate the true variance, resulting in approximate or below 90% coverage, when $m=30$ and covariate-adjusted working models (W1)-(W2) are used. In contrast, the MRS estimators, at worst, provide conservative coverage under comparable scenarios. Third, covariate adjustment can improve efficiency regardless of sample size and choice of estimation method. To summarize, under non-informative cluster size, the MRS method consistently performs no worse than the Coef method with a continuous outcome, although the latter has little bias despite model misspecification. This simulation experiment paves the way for Section 4.2 where we modify the data generating mechanism to induce informative cluster size with a continuous outcome.
@@ 4.2 Informative cluster size and a continuous outcome
Table 2: Simulation results under a data generating process with informative cluster size and a continuous outcome. LM: linear model based on cluster-level means; LMM: linear mixed model based on individual-level data; GEE-exch: marginal model fitted with GEE and an exchangeable working correlation structure; GEE-ind: marginal model fitted with GEE and working independence. Each empirical coverage probability is reported together with its Monte Carlo 95% confidence limits (95% CL). Additional Monte Carlo standard errors for performance metric estimates are presented in Web Table 2.
$m$ Covariate Outcome model Method BIAS MCSD COV BIAS MCSD COV
adjustment (AESE) (95% CL) (AESE) (95% CL)
@@
$30$ $\times$ (W1) LM Coef -0.3 2.16 (2.12) 93.6 (92.2, 95.0) -27.6 2.16 (2.12) 78.5 (76.0, 81.0)
MRS -0.3 2.16 (2.25) 94.5 (93.2, 95.8) -2.1 2.77 (2.88) 93.7 (92.3, 95.1)
(W2) LMM Coef -0.2 2.16 (2.12) 93.6 (92.2, 95.0) -27.6 2.16 (2.12) 78.5 (76.0, 81.0)
MRS -0.3 2.16 (2.25) 94.5 (93.2, 95.8) -2.1 2.77 (2.88) 93.7 (92.3, 95.1)
(W3) GEE-exch Coef -0.3 2.16 (2.28) 94.6 (93.3, 95.9) -27.6 2.16 (2.28) 81.7 (79.3, 84.1)
MRS -0.3 2.16 (2.25) 94.5 (93.2, 95.8) -2.1 2.77 (2.88) 93.7 (92.3, 95.1)
(W4) GEE-ind Coef 35.0 2.80 (2.99) 92.7 (91.3, 94.1) -2.0 2.80 (2.99) 93.8 (92.5, 95.1)
MRS -0.3 2.18 (2.31) 94.5 (93.2, 95.8) -2.0 2.80 (2.94) 93.4 (92.0, 94.8)
$\checkmark\checkmark$ (W1) LM Coef 0.1 1.18 (1.04) 90.7 (88.9, 92.5) -27.3 1.18 (1.04) 47.9 (44.8, 51.0)
MRS 0.1 1.18 (1.55) 97.4 (96.3, 98.5) -7.3 1.28 (1.71) 94.5 (93.2, 95.8)
(W2) LMM Coef 0.2 1.18 (1.04) 90.7 (88.9, 92.5) -27.3 1.18 (1.04) 48.0 (44.9, 51.1)
MRS 0.1 1.18 (1.43) 96.5 (95.4, 97.6) -7.3 1.28 (1.58) 93.6 (92.3, 94.9)
(W3) GEE-exch Coef 0.1 1.18 (1.46) 96.8 (95.8, 97.8) -27.3 1.18 (1.46) 68.7 (66.0, 71.4)
MRS 0.1 1.18 (1.43) 96.5 (95.4, 97.6) -7.3 1.28 (1.58) 93.6 (92.3, 94.9)
(W4) GEE-ind Coef 29.9 1.31 (1.66) 83.8 (81.6, 86.0) -5.7 1.31 (1.66) 94.7 (93.5, 95.9)
MRS 1.6 1.20 (1.49) 97.0 (96.0, 98.0) -5.7 1.31 (1.63) 94.1 (92.8, 95.4)
@@
$100$ $\times$ (W1) LM Coef -0.2 0.72 (0.74) 95.9 (94.8, 97.0) -28.4 0.72 (0.74) 34.6 (31.6, 37.6)
MRS -0.2 0.72 (0.75) 96.3 (95.3, 97.3) -1.1 0.85 (0.89) 95.2 (94.0, 96.4)
(W2) LMM Coef 0.0 0.72 (0.74) 95.8 (94.7, 96.9) -28.3 0.72 (0.74) 35.3 (32.3, 38.3)
MRS -0.2 0.72 (0.75) 96.3 (95.3, 97.3) -1.1 0.85 (0.89) 95.2 (94.0, 96.4)
(W3) GEE-exch Coef -0.3 0.72 (0.75) 96.3 (95.3, 97.3) -28.5 0.72 (0.75) 35.9 (32.9, 38.9)
MRS -0.2 0.72 (0.75) 96.3 (95.3, 97.3) -1.1 0.85 (0.89) 95.2 (94.0, 96.4)
(W4) GEE-ind Coef 37.9 0.83 (0.87) 51.2 (48.1, 54.3) -1.0 0.83 (0.87) 95.5 (94.4, 96.6)
MRS -0.2 0.74 (0.76) 96.2 (95.2, 97.2) -1.0 0.83 (0.86) 95.4 (94.3, 96.5)
$\checkmark\checkmark$ (W1) LM Coef 0.2 0.47 (0.44) 92.9 (91.4, 94.4) -28.1 0.47 (0.44) 3.7 (2.8, 4.6)
MRS 0.2 0.47 (0.49) 95.0 (93.9, 96.1) -2.1 0.52 (0.54) 93.8 (92.6, 95.0)
(W2) LMM Coef 0.7 0.47 (0.44) 93.1 (91.6, 94.6) -27.8 0.47 (0.44) 3.7 (2.8, 4.6)
MRS 0.2 0.47 (0.48) 94.9 (93.8, 96.0) -2.1 0.52 (0.52) 93.4 (92.2, 94.6)
(W3) GEE-exch Coef 0.3 0.47 (0.48) 94.9 (93.8, 96.0) -28.0 0.47 (0.48) 4.9 (3.8, 6.0)
MRS 0.2 0.47 (0.48) 94.9 (93.8, 96.0) -2.1 0.52 (0.52) 93.4 (92.2, 94.6)
(W4) GEE-ind Coef 37.1 0.51 (0.53) 10.8 (9.3, 12.3) -1.6 0.51 (0.53) 93.6 (92.4, 94.8)
MRS 0.6 0.47 (0.49) 94.7 (93.6, 95.8) -1.6 0.51 (0.52) 93.4 (92.2, 94.6)
@@
Table 2 presents the simulation results when the exact same 4 working models (with and without linear covariate adjustment) in Section 4.1 are fitted; the presentation of Table 2 follows Table 1 to enable direct comparisons. It is apparent that, under informative cluster size, the performance of the Coef estimator critically depends on the choice of the outcome regression model. In particular, regardless of covariate adjustment, the Coef estimator from the independence GEE (W4) is biased for estimating cluster-ATE, while the Coef estimators from the remaining models, (W1)-(W3), are biased for estimating individual-ATE. This bias then becomes the primary source for under-coverage. Such biases have been pointed out from previous studies,11, 12 and arise because the fitted regression model does not employ the right weighting scheme that matches the marginal estimands $\Delta_{C}$ or $\Delta_{I}$ (also see Kahan et al.5 and Lee et al.lee2025should for detailed explanation about why Coef from LMM could be biased under informative cluster size). In sharp contrast, the relative bias is negligible and the empirical coverage probabilities are close to $95\%$ for MRS estimators, regardless of the choice of outcome regression models and covariate adjustment. Finally, covariate adjustment under MRS also leads to no worse finite-sample efficiency compared to no covariate adjustment. In summary, although Coef and MRS perform relatively similarly under non-informative cluster size, the findings in Section 4.2 suggest that (1) the performance of Coef highly depends on the choice of outcome model and Coef may target neither cluster-ATE or individual-ATE; (2) even with a given outcome model whose Coef estimator is biased for the estimand of interest, applying our MRS procedure can fix the bias.
For readers interested in isolating specific sources of model misspecification on the results patterns, we have considered a simpler setup where informative cluster size is introduced only through the treatment effect (rather than complex covariate dependence and ICC heterogeneity). The detailed simulation results are reported in Web Appendix Section D, and the overall trends are similar to those observed in Table 2. This reinforces the robustness of our main findings.